The equations x-y=4, x^2+4xy+y^2=0 represent sides of 
1.equilateral triangle 2.right angled triangle 3.isosceles triangle

Grade:

2 Answers

Manda Snehashish Reddy
5 years ago
Please find your answer below
REGARDS, Ashish
 
NOTE: As you are as aspirant, I’d rather suggest you to solve the equations yourself. Do it yourself like many others to improve your speed and accuracy.
Manda Snehashish Reddy
5 years ago
x-y = 4
x2+4xy+y2=0
Simplifying the above equation
x2+4xy+y2+3y2=3y2
(x+2y)2 =3y2
y =(31/2 -2)x or y = -(31/2+2)x
These two lines intersect at (0,0)
x-y=4 intersects with x2+4xy+y2=0 at (0.845,-3.155) and (3.155,-0.845)

Lengths of the sides are 3.266,3.266,3.266

Therefore the triangle is an equilateral triangle.